A golf ball with a mass of 0.05 kg is truck with a club. The duration of the collision between the golf ball and the club is a fraction of a second. Assume the ball leaves the club face with a velocity of +44 m/s. Find the magnitude of the impulse due to the collision.
Answer #1

Impulse = Change in momentum

=M×(v2-v1)

= 0.05×(44 - 0) = 2.2 kg m/s